# A Survey of Self-Evolving Agents: On Path to Artificial Super Intelligence

### v3

@@ 2.2 Relationships with Other Works
Table 1 summarizes the key distinctions between self-evolving agents and other paradigms (including curriculum learning, lifelong learning, model editing, and unlearning). Unlike these existing paradigms that primarily focus on updating model parameters, self-evolving agents expand the scope of updating targets to include non-parametric components, such as context (prompts and memory) and toolset. This expanded space provides greater flexibility, enabling self-evolving agents to operate effectively in sequential task settings and adapt at test time. More crucially, self-evolving agents uniquely demonstrate the ability of active exploration (e.g., searching for open-source tools online qiu2025alita ), structural modification of their own topology (e.g., iteratively modifying the workflow zhang2024aflow or code zhang2025darwin ), and self-reflection and self-evaluation capabilities (e.g., providing verbal feedback using an internal evaluator LLM shinn2023reflexion ), which are absent in previous paradigms.
We provide a brief introduction to each paradigm below, highlighting the differences among these paradigms, as well as the differences with self-evolving agents.
@@
Curriculum learning is a training strategy for AI models in which data are presented in order of increasing difficulty bengio2009curriculum ; wang2021survey . This strategy resembles human curricula where concepts are introduced progressively from simple to complex. Curriculum learning has been widely adopted across diverse domains, including computer vision guo2018curriculumnet ; jiang2014easy ; liu2023towards , natural language processing platanios2019competence ; tay2019simple , speech recognition braun2017curriculum ; lotfian2019curriculum , etc. Recently, several curriculum learning-based methods have been proposed to fine-tune LLMs during the post-training phase wang2025dump ; zhang2025preference ; parashar2025curriculum ; zhang2025learning ; li2025curriculum . The framework for curriculum learning generally comprises two key components: a difficulty measurer that quantifies the difficulty level of each training data point, and a training scheduler that reorganizes the order of data points received by the model according to the difficulty level. Unlike curriculum learning, which operates on a static dataset, self-evolving agents aim to handle sequential tasks in dynamic environments. Additionally, curriculum learning updates only model parameters, whereas self-evolving agents are able to adjust non-parametric components like memory and tools.
@@
Lifelong learning refers to the ability of AI models to continuously and adaptively learn when exposed to new tasks and environments, while retaining previously acquired knowledge and abilities. This learning paradigm, also known as continual learning or incremental learning, is crucial for AI models to operate in dynamic and complex environments wang2024comprehensive ; zheng2025lifelong ; parisi2019continual ; shi2024continual ; yang2025recent ; zhou2024continual . The primary goal of lifelong learning for AI models is to achieve a balance between preserving existing knowledge (stability) and acquiring new knowledge (plasticity) when exposed to new data or tasks mccloskey1989catastrophic ; zheng2025lifelong ; ratcliff1990connectionist ; rolnick2019experience . Though it shares the sequential task setting with self-evolving agents, lifelong learning differs in two ways: (1) like curriculum learning, lifelong learning typically updates only model parameters, lacking the ability to modify non-parametric components; (2) lifelong learning primarily acquires knowledge passively through external feedback or manual guidance, whereas self-evolving agents actively explore their environment and may incorporate internal reflection or self-evaluation mechanisms.
@@
Model editing and unlearning aim to efficiently and precisely modify specific knowledge in AI models while preserving irrelevant knowledge and avoiding full retraining wang2024knowledge ; wang2025comprehensive ; zhang2024comprehensive ; wang2025comprehensive ; nguyen2022survey ; geng2025comprehensive . A canonical application of model editing is to perform efficient and precise localized factual updates (e.g., modifying the answer to “2021 Olympics host city” from "Tokyo" to "Paris"). Early methods focused on triples of atomic knowledge and later expanded into various trustworthy-related tasks fang2025alphaedit ; huang2025model . Recent studies also propose lifelong model editing chen2024lifelong that sequentially performs model editing. For model unlearning, early efforts mainly focus on the removal of privacy-related information chen2021machine . With the rapid development of LLMs, model unlearning is also used to enhance LLMs’ safety safe_unlearning ; li2024wmdp ; circuit_breaker ; lu2025x . Compared to lifelong learning, model editing shares an aligned objective: both aim to acquire new knowledge or capabilities while mitigating catastrophic forgetting. However, lifelong learning typically relies on extensive gradient-based fine-tuning across all model parameters, whereas model editing often modifies only a small subset of parameters in a targeted manner. Compared to self-evolving agents, model editing (1) cannot modify non-parametric components such as memory or tools, and (2) relies on a pre-defined pipeline from the algorithm designer, whereas self-evolving agents can spontaneously employ more diverse and flexible strategies based on the observation of the environment or internal feedback signals.
Table 1: Comparison between self-evolving agents and other renowned paradigms, including curriculum learning, lifelong learning, model editing and unlearning.
Paradigm Evolving Evolving Dynamic Test-time Active Structural Self-reflect
@@ 3 What to Evolve?
The transition from pre-programmed, static systems to genuinely autonomous intelligent agents hinges on one critical capability: self-evolution. This capability for continuous improvement is not monolithic; instead, it manifests across various facets of an agent’s being. Therefore, the first key aspect of a self-evolving agent is identifying the evolving components – which components in the agent system $\Pi=(\Gamma,\{\psi_{i}\},\{C_{i}\},\{\mathcal{W}_{i}\})$ can be explicitly changed over time?
Following the formulation in Section 2.1, we can decompose the agent system into four fundamental, evolvable pillars. Our investigation starts at the agent’s cognitive core, namely the Models $\{\psi_{i}\}$, and examines how the fundamental parameters that govern its reasoning and behavior are continuously updated through its own experienceszhou2025selfchallenginglanguagemodelagents ; wang2025ragen . We then consider the Context $\{C_{i}\}$ that shapes its actions, exploring the evolution of both the instructionsxiang2025self ; khattab2023dspy it follows and the long-term memorychhikara2025mem0 ; wang2024agent it draws upon to stay informed and adapt. From this internal foundation, we shift to the agent’s external capabilities, analyzing how it evolves its set of Tools $\{\mathcal{W}_{i}\}$ by autonomously creatingqiu2025alita , masteringfromexplorationtomastery , and managing new skillswang2025toolgen to overcome its innate limitations. Finally, we scale our perspective to the Agentic System itself, investigating how the agentic system’s architecturehu2024automated ; zhang2024aflow and collaborative structureswan2025remalearningmetathinkllms are dynamically optimized to enhance overall performance and efficiency. We present a subset of these evolving dimensions in Table 2.
@@ 3.1 Models
Models constitute the central substrate of intelligent agents, directly determining their reasoning, planning, and decision-making behaviors. The ability of these models to evolve by continually adapting their internal parameters and expanding their functional capabilities is essential for the development of autonomous, general-purpose agents. Unlike static systems that rely heavily on human-annotated datasets and fixed training regimes, self-evolving models can improve through interaction, self-supervised data generation, and dynamic learning loops, thereby achieving greater efficiency, adaptability, and scalability. In detail, we outline the principal axes along which model evolution unfolds. These include learning from self-generated supervision to refine model weights, evolving through interaction with constructed or external environments. Together, these strategies represent a shift from passive learning paradigms toward active, continual, and self-directed improvement.
@@
SCA zhou2025selfchallenginglanguagemodelagents $\bullet$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$
RAGEN wang2025ragen $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
AgentGen  hu2024agentgen $\circ$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\circ$
Promptbreeder fernando2023promptbreeder $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
Expel zhao2024expel $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Agent Workflow Memory wang2024agent $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$
Mem0 chhikara2025mem0 $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
MAS-Zero ke2025maszero $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$
Multi-Agent Design zhou2025maasprompt $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$
SPO xiang2025self $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Alita qiu2025alita $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$
TextGrad yellamraju2024textgrad $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\circ$
DGM zhang2025darwin $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
AlphaEvolve novikov2025alphaevolve $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\bullet$ $\circ$
ADAS hu2024automated $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\bullet$
AFlow zhang2024aflow $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\bullet$ $\bullet$
ReMA wan2025remalearningmetathinkllms $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$
SkillWeaver zheng2025skillweaver $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$
LearnAct zhao2024learnact $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$
DRAFT fromexplorationtomastery $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$
ToolGen wang2025toolgen $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$
CRAFT yuan2024craft $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$
CREATOR qian2023creator $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$
Voyager wang2023voyager $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\bullet$
@@
A self-evolving agent can refine its parameters to perform better on targeted tasks. Traditional methods of data collection for training agents on tool-use benchmarks are costly and often yield limited coverage, while purely synthetic data-generation pipelines typically suffer from inadequate quality. Consequently, recent studies emphasize enabling agents to autonomously generate data to improve their own model weights. One representative approach is the Self-Challenging Agent (SCA)zhou2025selfchallenginglanguagemodelagents , where a language model alternates roles between a challenger generating executable Code-as-Task problems and an executor solving them. The model then fine-tunes its parameters using trajectories derived from successful solutions, resulting in significant performance gains on complex, multi-step tasks. Similarly, the Self-Rewarding Self-Improving frameworksimonds2025selfrewardingselfimproving implements an internal self-judging mechanism, allowing the model to autonomously generate problems, solve them, and assess its performance, thus producing self-contained fine-tuning data without external annotations. This method demonstrated notable improvements, particularly in complex reasoning tasks. Beyond task creation, another promising research direction involves leveraging interaction feedback directly for parameter updates. For instance, SELF lu2023self , SCoRe kumar2024training , and PAG jiang2025pag interpret execution traces or natural-language critiques as reward signals within an online Supervised Fine-Tuning (SFT) combined with Reinforcement Learning (RL) framework, enabling continuous policy improvement. TextGrad yellamraju2024textgrad further extends this concept by treating unstructured textual feedback as a differentiable training signal capable of directly influencing both prompt design and model parameters. Additionally, AutoRule wang2025autorule converts language-model reasoning traces and preference feedback into explicit rule-based training rewards, enhancing the quality of model outputs through structured reward signals. Collectively, these advancements chart a clear trajectory—from agents autonomously crafting their training tasks to directly refining their parameters based on execution feedback, highlighting the capacity of models to evolve continuously by learning from the data they produce.
@@
Agents can evolve not only by adjusting their internal parameters but also by actively interacting with or even constructing their environments, capturing experiences, and transforming them into learning signals that drive iterative improvement. This environmental loop provides agents with the complexity and diversity required for scalable self-adaptation. The Self-Challenging Agent (SCA)zhou2025selfchallenginglanguagemodelagents exemplifies this dynamic at the task level, where the agent autonomously generates novel Code-as-Task problems, executes them, and then filters successful trajectories for retraining itself. AgentGenhu2024agentgen extends this concept to full-environment generation, synthesizing diverse simulation worlds (in PDDL or Gym-style formats) derived from an initial corpus. It implements a bidirectional evolution loop that progressively adjusts task difficulty, enabling the agent to continuously grow within a dynamically structured curriculum. Reflexion shinn2023reflexion complements this by introducing self-reflective mechanisms, where agents iteratively record natural-language critiques of their previous actions, guiding future behavior to avoid recurring mistakes. Additionally, AdaPlanner sun2023adaplanner introduces closed-loop adaptive planning, allowing agents to refine their strategies on-the-fly based on environmental feedback, effectively reshaping action sequences in response to immediate outcomes. Similarly, Self-Refine madaan2023selfrefine employs an iterative refinement loop in which the agent repeatedly critiques and revises its initial outputs, significantly improving task accuracy without explicit retraining. SICA (Self-Improving Coding Agent) robeyns2025selfimprovingcodingagent further pushes the boundary by enabling agents to autonomously edit their underlying code and tools, iteratively enhancing their core reasoning abilities through direct self-modification. From a reinforcement learning perspective, frameworks such as RAGEN wang2025ragen and DYSTIL wang2025dystildynamicstrategyinduction conceptualize multi-step tool-use tasks as Markov Decision Processes, optimizing agent policies through rich environmental rewards and strategy induction loops. RAGEN leverages dense feedback from the environment to iteratively fine-tune action policies, while DYSTIL utilizes high-level strategy advice generated by language models to progressively internalize complex decision-making skills into reinforcement learning agents. Collectively, these approaches highlight a compelling paradigm where self-evolving agents not only leverage self-generated data but actively reshape their environments and internal mechanisms to fuel ongoing learning. Such dynamic interaction loops point toward autonomous, open-ended improvement cycles deeply grounded in experiential adaptation.

### v4

@@ 2.2 Relationships with Other Works
Table 1 summarizes the key distinctions between self-evolving agents and other paradigms (including curriculum learning, lifelong learning, model editing, and unlearning). We provide a brief introduction to each paradigm below, highlighting the differences among these paradigms, as well as the differences with self-evolving agents.
@@
Curriculum learning is a training strategy in which data are presented in order of increasing difficulty (Bengio et al., 2009; Wang et al., 2021). This strategy resembles human curricula where concepts are introduced progressively from simple to complex. Curriculum learning has been widely adopted across diverse domains, including computer vision (Guo et al., 2018; Jiang et al., 2014; Liu et al., 2023a), natural language processing (Platanios et al., 2019; Tay et al., 2019), speech recognition (Braun et al., 2017; Lotfian and Busso, 2019), etc. Recently, several curriculum learning-based methods have been proposed to fine-tune LLMs during the post-training phase (Wang et al., 2025o; Zhang et al., 2025o; Parashar et al., 2025; Zhang et al., 2025a; Li et al., 2025b). The framework for curriculum learning generally comprises two key components: a difficulty measurer that quantifies the difficulty level of each training data point, and a training scheduler that reorganizes the order of data points received by the model according to the difficulty level. Unlike curriculum learning, which operates on a static dataset, self-evolving agents aim to handle sequential tasks in dynamic environments. Additionally, curriculum learning updates only model parameters, whereas self-evolving agents are able to adjust non-parametric components like memory and tools.
@@
Lifelong learning refers to the ability of AI models to continuously and adaptively learn when exposed to new tasks and environments, while retaining previously acquired knowledge and abilities. This learning paradigm, also known as continual learning or incremental learning, is crucial for AI models to operate in dynamic and complex environments (Wang et al., 2024c; Zheng et al., 2025c; Parisi et al., 2019; Shi et al., 2024; Yang et al., 2025d; Zhou et al., 2024a). The primary goal of lifelong learning for AI models is to achieve a balance between preserving existing knowledge (stability) and acquiring new knowledge (plasticity) when exposed to new data or tasks (McCloskey and Cohen, 1989; Zheng et al., 2025c; Ratcliff, 1990; Rolnick et al., 2019). Though it shares the sequential task setting with self-evolving agents, lifelong learning differs in two fundamental ways: (1) Memory functionality and usage timing: While continual learning methods extensively employ memory mechanisms (e.g., experience replay buffers (Rolnick et al., 2019), episodic memory (Lopez-Paz and Ranzato, 2017)) to mitigate catastrophic forgetting, these mechanisms primarily serve as training-time tools for parameter optimization through gradient computation. In contrast, self-evolving agents leverage runtime context (prompts, working memory, conversation history) that directly influences action generation at test-time without requiring parameter updates. The distinction lies not in the presence of non-parametric components, but in their functional role: training-time replay vs. test-time state adaptation. (2) Learning initiative: Lifelong learning primarily acquires knowledge passively through externally provided task sequences, whereas self-evolving agents actively explore their environment and incorporate internal reflection or self-evaluation mechanisms to guide their own learning trajectory. Recent self-improving LLM methods (Huang et al., 2022; Yuan et al., 2024c), which iteratively refine models through self-generated data and self-critique, can be viewed as instances of lifelong learning focused on model-centric improvement. Self-evolving agents extend beyond this paradigm to encompass system-wide evolution including tool acquisition, architectural reconfiguration, and environmental exploration.
@@
Model editing and unlearning aim to efficiently and precisely modify specific knowledge in AI models while preserving irrelevant knowledge and avoiding full retraining (Wang et al., 2024f; 2025h; Zhang et al., 2024d; Wang et al., 2025h; Nguyen et al., 2022; Geng et al., 2025a). A canonical application of model editing is to perform efficient and precise localized factual updates (e.g., modifying the answer to "2021 Olympics host city" from "Tokyo" to "Paris"). Early methods focused on triples of atomic knowledge and later expanded into various trustworthy-related tasks (Fang et al., 2025a; Huang et al., 2025a). Recent studies also propose lifelong model editing(Chen et al., 2024c) that sequentially performs model editing. For model unlearning, early efforts mainly focus on the removal of privacy-related information (Chen et al., 2021). With the rapid development of LLMs, model unlearning is also used to enhance LLMs’ safety (Zhang et al., 2024j; Li et al., 2024c; Zou et al., 2024; Lu et al., 2025). Compared to lifelong learning, model editing shares an aligned objective: both aim to acquire new knowledge or capabilities while mitigating catastrophic forgetting. However, lifelong learning typically relies on extensive gradient-based fine-tuning across all model parameters, whereas model editing often modifies only a small subset of parameters in a targeted manner. Compared to self-evolving agents, model editing (1) cannot modify non-parametric components such as memory or tools, and (2) relies on a pre-defined pipeline from the algorithm designer, whereas self-evolving agents can spontaneously employ more diverse and flexible strategies based on the observation of the environment or internal feedback signals.
Table 1: Comparison between self-evolving agents and other renowned paradigms
Paradigm Runtime Evolving Dynamic Test-time Active Structural Self-reflect
@@
Positioning Self-Evolving Agents
To clarify the relationships among these paradigms and to motivate the role of self-evolving agents, we examine them through two complementary perspectives: a problem-setting lens and a solution-paradigm lens. This distinction clarifies the basis of each paradigm - whether it emerges from constraints and challenges inherent to the learning setting, or from methodological proposals for how the model or agent itself can be updated.
  • Problem-setting view. Curriculum learning and lifelong learning arise from concrete learning problems. Curriculum learning addresses how to structure training examples of varying difficulty so a model can handle complex samples more effectively; lifelong learning focuses on acquiring new abilities over time while mitigating catastrophic forgetting. These paradigms are therefore driven by the problems they aim to solve and primarily specify how experience is organized for the learner, rather than how the agent itself may adapt beyond parameter updates.
  • Solution-paradigm view. Model editing and self-evolving agents, in contrast, originate as solutions: they propose mechanisms for updating or modifying a system. Model editing provides targeted procedures—typically localized parameter adjustments—to correct or insert knowledge. Self-evolving agents generalize this idea by treating adaptation as a first-class capability, allowing not only parameter updates but also changes to runtime context, memory, tools, and workflow structures, driven by the agent’s own trajectories and feedback signals.
Viewed through this two-lens framework, curriculum and lifelong learning are anchored in the nature of the learning problems they address, whereas model editing and self-evolving agents are defined by the methods they provide for effecting change. Self-evolving agents thus represent a system-level solution paradigm: they include parameter-level editing as one update pathway while enabling broader, persistent, and interaction-driven evolution across multiple components of an agent.
@@ 3 What to Evolve?
A self-evolving agent differs from a static agent not by what components it contains, but by which internal states can be autonomously modified based on its own trajectories, reflections, and feedback signals. Thus, the key question of this section is to identify the evolutionary loci within an agent system $\Pi=(\Gamma,\{\psi_{i}\},\{C_{i}\},\{\mathcal{W}_{i}\})$—the parts of the system whose states can be rewritten in an experience-driven and persistent manner, enabling cumulative self-improvement.
Following the formulation in Section 2.1, these evolutionary loci align with four major pillars of an agent system. Our investigation starts at the agent’s cognitive core, namely the Models $\{\psi_{i}\}$, whose parameters can be continuously updated through self-generated supervision, execution traces, or environmental feedback (Zhou et al., 2025e; Wang et al., 2025p). We then consider the Context $\{C_{i}\}$ –including instructions (Xiang et al., 2025; Khattab et al., 2023) and long-term memory (Chhikara et al., 2025; Wang et al., 2024j) –which evolves as agents reflect, store, and retrieve experience in ways that shape future decision-making. From this internal foundation, we examine the evolution of Tools $\{\mathcal{W}_{i}\}$, where agents autonomously create (Qiu et al., 2025b), refine (Qu et al., 2025), and managing (Wang et al., 2025j) executable skills based on verifiable interaction signals Finally, we scale to the Agentic Architecture, where the system’s architecture (Hu et al., 2024c; Zhang et al., 2024c) and collaborative structures (Wan et al., 2025) are optimized over time, enabling structural adaptation beyond individual components. We present representative examples of these evolutionary loci in Table 2.
@@ 3.1 Models
Models constitute a primary locus of self-evolution, as their parameters can be autonomously rewritten based on the agent’s own trajectories, reflections, and interaction outcomes. The ability of these models to evolve by continually adapting their internal parameters and expanding their functional capabilities is essential for the development of autonomous, general-purpose agents. Unlike static systems that rely heavily on human-annotated datasets and fixed training regimes, self-evolving models can improve through interaction, self-supervised data generation, and dynamic learning loops, thereby achieving greater efficiency, adaptability, and scalability.
In detail, we outline the principal axes along which model evolution unfolds. These include learning from self-generated supervision to refine model weights, evolving through interaction with constructed or external environments, and integrating feedback signals that directly reshape future reasoning behaviors. Together, these strategies represent a shift from passive learning paradigms toward active self-improvement.
@@
SCA(Zhou et al., 2025e) $\bullet$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$
RAGEN(Wang et al., 2025p) $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
AgentGen (Hu et al., 2024b) $\circ$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\circ$
Promptbreeder(Fernando et al., 2023) $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
Expel(Zhao et al., 2024a) $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Agent Workflow Memory(Wang et al., 2024j) $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$
Mem0(Chhikara et al., 2025) $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
MAS-Zero(Ke et al., 2025) $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$
Multi-Agent Design(Zhou et al., 2025a) $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$
SPO(Xiang et al., 2025) $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Alita(Qiu et al., 2025b) $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$
TextGrad(Yellamraju and others, 2024) $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\circ$
DGM(Zhang et al., 2025h) $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
AlphaEvolve(Novikov et al., 2025) $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\bullet$ $\circ$
ADAS(Hu et al., 2024c) $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\bullet$
AFlow(Zhang et al., 2024c) $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\bullet$ $\bullet$
ReMA(Wan et al., 2025) $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$
SkillWeaver(Zheng et al., 2025a) $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$
LearnAct(Zhao et al., 2024b) $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$
DRAFT(Qu et al., 2025) $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$
ToolGen(Wang et al., 2025j) $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$
CRAFT(Yuan et al., 2024a) $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$
CREATOR(Qian et al., 2023b) $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$
Voyager(Wang et al., 2023a) $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\bullet$
@@
A self-evolving agent can refine its parameters to perform better on targeted tasks. Traditional methods of data collection for training agents on tool-use benchmarks are costly and often yield limited coverage, while purely synthetic data-generation pipelines typically suffer from inadequate quality. Consequently, recent studies emphasize enabling agents to autonomously generate data to improve their own model weights. One representative approach is the Self-Challenging Agent (SCA)(Zhou et al., 2025e), where a language model alternates roles between a challenger generating executable Code-as-Task problems and an executor solving them. The model then fine-tunes its parameters using trajectories derived from successful solutions, resulting in significant performance gains on complex, multi-step tasks. Similarly, the Self-Rewarding Self-Improving framework(Simonds et al., 2025) implements an internal self-judging mechanism, allowing the model to autonomously generate problems, solve them, and assess its performance, thus producing self-contained fine-tuning data without external annotations. This method demonstrated notable improvements, particularly in complex reasoning tasks. Beyond task creation, another promising research direction involves leveraging interaction feedback directly for parameter updates. For instance, SELF(Lu et al., 2023), SCoRe(Kumar et al., 2024), and PAG(Jiang et al., 2025b) interpret execution traces or natural-language critiques as reward signals within an online Supervised Fine-Tuning (SFT) combined with Reinforcement Learning (RL) framework, enabling continuous policy improvement. TextGrad(Yellamraju and others, 2024) further extends this concept by treating unstructured textual feedback as a differentiable training signal capable of directly influencing both prompt design and model parameters. Additionally, AutoRule(Wang and Xiong, 2025) converts language-model reasoning traces and preference feedback into explicit rule-based training rewards, enhancing the quality of model outputs through structured reward signals. Collectively, these advancements chart a clear trajectory—from agents autonomously crafting their training tasks to directly refining their parameters based on execution feedback, highlighting the capacity of models to evolve continuously by learning from the data they produce.
@@
Agents can evolve not only by adjusting their internal parameters but also by actively interacting with or even constructing their environments, capturing experiences, and transforming them into learning signals that drive iterative improvement. This environmental loop provides agents with the complexity and diversity required for scalable self-adaptation. The Self-Challenging Agent (SCA)(Zhou et al., 2025e) exemplifies this dynamic at the task level, where the agent autonomously generates novel Code-as-Task problems, executes them, and then filters successful trajectories for retraining itself. AgentGen(Hu et al., 2024b) extends this concept to full-environment generation, synthesizing diverse simulation worlds (in PDDL or Gym-style formats) derived from an initial corpus. It implements a bidirectional evolution loop that progressively adjusts task difficulty, enabling the agent to continuously grow within a dynamically structured curriculum. Reflexion(Shinn et al., 2023) complements this by introducing self-reflective mechanisms, where agents iteratively record natural-language critiques of their previous actions, guiding future behavior to avoid recurring mistakes. Additionally, AdaPlanner(Sun et al., 2023) introduces closed-loop adaptive planning, allowing agents to refine their strategies on-the-fly based on environmental feedback, effectively reshaping action sequences in response to immediate outcomes. Similarly, Self-Refine(Madaan et al., 2023b) employs an iterative refinement loop in which the agent repeatedly critiques and revises its initial outputs, significantly improving task accuracy without explicit retraining. SICA (Self-Improving Coding Agent)(Robeyns et al., 2025b) further pushes the boundary by enabling agents to autonomously edit their underlying code and tools, iteratively enhancing their core reasoning abilities through direct self-modification. From a reinforcement learning perspective, frameworks such as RAGEN(Wang et al., 2025p) and DYSTIL(Wang et al., 2025c) conceptualize multi-step tool-use tasks as Markov Decision Processes, optimizing agent policies through rich environmental rewards and strategy induction loops. RAGEN leverages dense feedback from the environment to iteratively fine-tune action policies, while DYSTIL utilizes high-level strategy advice generated by language models to progressively internalize complex decision-making skills into reinforcement learning agents. Collectively, these approaches highlight a compelling paradigm where self-evolving agents not only leverage self-generated data but actively reshape their environments and internal mechanisms to fuel ongoing learning. Such dynamic interaction loops point toward autonomous, open-ended improvement cycles deeply grounded in experiential adaptation.